A kangaroo is capable of jumping to a height of 2.62 m. determine the takeoff speed of the kangaroo.
1
Expert's answer
2012-01-31T12:40:18-0500
Due to energy conservation law

Mgh=Mv2/2
where M-mass of kangaroo, v- take off speedHence v=sqrt(2gh)=sqrt(2*10*2.62)=7.2 m/s